OpenStudy (anonymous):

h(b1+b2) divided by 2

OpenStudy (anonymous):

That is the formula for the area of a trapezoid

OpenStudy (anonymous):

how do you do it

OpenStudy (anonymous):

what do you mean? you need numbers

hartnn (hartnn):

@jadasmi Hi, \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\) Is your question complete? can u please verify. solve for what?

OpenStudy (anonymous):

its h(b1+b2) divided by 2

hartnn (hartnn):

ok, i think you need to solve for some variable like h, or b1 or b2 can u tell me which topic does this question belong? or can u post a screenshot of the question ? so that we can help you better

OpenStudy (anonymous):

ok im solveing for b1 the formula is a=h(b1 + b2) divided by 2

hartnn (hartnn):

see, now this is correct question :) to isolate b1, we need to bring everything that is with b1 to other side. 2 is divided, so now can u multiply 2 on both sides and tell me what u get ?

OpenStudy (anonymous):

o divide b1 to bot sides

hartnn (hartnn):

nopes, a= h(b1+b2) / 2 so 2a = h(b1+b2) <---multiplying 2 on both sides did u get this ?

OpenStudy (anonymous):

so i divide 2 to h

OpenStudy (anonymous):

what do i divided to

hartnn (hartnn):

since h is multiplied, to get rid of h, you divide h on both sides. ok ? so 2a/h = b1+ b2 did u get this ?

OpenStudy (anonymous):

yess

hartnn (hartnn):

now since b2 is ADDED to b1, to get rid of it, we SUBTRACT b2 from both sides, what u get ?

OpenStudy (anonymous):

i got 0

hartnn (hartnn):

lol :P 2a/h = b1+b2 2a/h - b2 = b1 so u get b1 = 2a/h -b2 got this ?

hartnn (hartnn):

that would be your final answer

OpenStudy (anonymous):

ok i got it

hartnn (hartnn):

great :)

OpenStudy (anonymous):

thank you